OpenStudy (anonymous):

Which number has the least absolute value? A. -100 B. -25 C. -12 D. 10 PLEASE EXPLAIN!!!

sammixboo (sammixboo):

Do you know what absolute value is?

Mehek (mehek14):

Absolute value means the distance from the number to 0 doesn't matter about the sign -100 = 100 -25 = 25 -12 = 12 10 = 10

sammixboo (sammixboo):

Let me give you some examples. The absolute value of -5 is 5. |-5| = 5 Also, the absolute value of 5 is 5, |5| = 5

sammixboo (sammixboo):

All you do is strip away the integer sign

OpenStudy (anonymous):

Ok so the answer would be C? @sammixboo

sammixboo (sammixboo):

Not quite. Almost, but no

OpenStudy (anonymous):

d

sammixboo (sammixboo):

The absolute value of -12 is 12, but look at 10, what is the absolute value of that? 10! so your answer is D, correct

OpenStudy (anonymous):

Ok thank you so much (:

sammixboo (sammixboo):

Absolute value of positive numbers are the same

OpenStudy (anonymous):

It is greatest because it is 100 away from zero. The rest are less than that.